OpenStudy (anonymous):

Which of the following represents the set of possible rational roots for the polynomial shown below? x^3+5x^2-8x-20=0 A.{1/2, 1,2, 5/2, 4, 5, 10, 20} B. {+/-1, +/-2, +/-4, +/-5, +/-10} C. {+/-1/2, +/-1, +/-2, +/-5/2, +/-4, +/-5, +/-10, +/-20} D. {+/-2/5, +/-1/2, +/-1, +/-2, +/-2/5, +/-1/5, +/-1/10}

OpenStudy (anonymous):

You want to write out the factors of 20 (the last number) first: 1, 2, 4, 5, 10, 20 Lets call these guys 'p' Then you want to write out the factors of the leading coefficient (the number in front of the highest power of x): 1 Lets call this guy 'q' Then you want to create all the possible fractions you can with: \[\pm \frac{p}{q}\] Since q is just 1, we are lucky, so the possible rational roots are: \[\pm \frac{1}{1}, \pm \frac{2}{1}, \pm \frac{4}{1}, \pm \frac{5}{1}. \pm \frac{10}{1}, \pm \frac{20}{1}\]

OpenStudy (anonymous):

This is called the Rational Roots Theorem. http://www.mathwords.com/r/rational_root_theorem.htm

OpenStudy (anonymous):

If you need anything explained in more detail just ask :)

OpenStudy (anonymous):

thank you :)